The lengthy of a rectangular garden is 2 yd greater than the width. The area is 80 yd squared. Find the length and width.
Answer by stanbon(75887)

The length of a rectangular garden is 2 yd greater than the width. The area is 80 yd squared. Find the length and width.
-------------
Let width be "x"
Then y = "x+2".
-------
Equation:
area = LW
80 yd^2 = x(x+2)
x^2+2x-80 = 0
-----
Positive solution:
x = 8 yd (width)
x+2 = 10 yd (length)
==================
Cheers,
Stan H.